### 18 Mar 2024 v2

# Fisher Mask Nodes for Language Model Merging

Abstract
Fine-tuning pre-trained models provides significant advantages in downstream performance. The ubiquitous nature of pre-trained models such as BERT and its derivatives in natural language processing has also led to a proliferation of task-specific fine-tuned models. As these models typically only perform one task well, additional training or ensembling is required in multi-task scenarios. The growing field of model merging provides a solution, dealing with the challenge of combining multiple task-specific models into a single multi-task model. In this study, we introduce a novel model merging method for Transformers, combining insights from previous work in Fisher-weighted averaging and the use of Fisher information in model pruning. Utilizing the Fisher information of mask nodes within the Transformer architecture, we devise a computationally efficient weighted-averaging scheme. Our method exhibits a regular and significant performance increase across various models in the BERT family, outperforming full-scale Fisher-weighted averaging in a fraction of the computational cost, with baseline performance improvements of up to +6.5 and a speedup of 57.4x in the biggest model. Our results prove the potential of our method in current multi-task learning environments and suggest its scalability and adaptability to new model architectures and learning scenarios.
Keywords: model merging, Fisher information, NLP, GLUE, transformers, BERT
[…]
languageresource
Fisher Mask Nodes for Language Model Merging
Thennal D K, Ganesh Nathan, Suchithra M S
Indian Institute of Information Technology Kottayam
Kerala, India
{thennal21bcs14, ganesh21bcs104, suchithra}@iiitkottayam.ac.in
Abstract content
1.   Introduction
In recent years, pre-trained models have become ubiquitous in natural language processing (NLP) (Qiu et al., 2020; Han et al., 2021; Zhou et al., 2023; Min et al., 2021). In particular, models inheriting the Transformer architecture (Vaswani et al., 2017) such as BERT and its variants (Devlin et al., 2019) are popular for a wide variety of tasks (Zhou et al., 2023; Min et al., 2021). By finetuning on downstream tasks, pre-trained models achieve improved performance on those specific tasks. However, individual fine-tuned models for each task necessitates significant overhead in multi-task scenarios, as multiple separate models need to be running at once (Zhang and Yang, 2021; Fifty et al., 2021; Yadav et al., 2023). Furthermore, individual fine-tuned models sacrifice generalization capabilities (Jin et al., 2023; Arpit et al., 2022; Yadav et al., 2023; Ilharco et al., 2023). Fine-tuning on multiple tasks at once is a possible solution, but is generally resources-intensive and costly (Yadav et al., 2023; Fifty et al., 2021).
Model merging is a promising field that can alleviate this problem. Model merging predominantly deals with the challenges of fusing multiple separate models—generally of the same architecture but with different parameters—into a single model. A growing body of literature addresses the restricted context of merging multiple task-specific models into a single multitask model (Yadav et al., 2023; Matena and Raffel, 2022; Ilharco et al., 2023; Jin et al., 2023). Fisher-weighted averaging (Matena and Raffel, 2022) is one such method that uses the diagonal approximation of the Fisher information of each parameter as weights in a weighted-averaging scheme. However, doing so requires access to the validation set that the model will be evaluated on, along with significant resource overhead, as gradients for all parameters are required to calculate the approximated Fisher information.
Fisher information has been used extensively in machine learning as a measure of parameter importance (Kwon et al., 2022; Liu et al., 2021; Kirkpatrick et al., 2017a; Pascanu and Bengio, 2014; Soen and Sun, 2021). In recent years, it has gained particular prominence in the field of model pruning (Liu et al., 2021; Kwon et al., 2022). Kwon et al. (2022) devised a pruning scheme for Transformer models utilizing Fisher information matrices. Masks are inserted in every feed-forward layer and attention head, and the approximated Fisher information of the masks is used as an indication of the necessity of the affected block of parameters. Low-importance parameters, indicated by this metric, are then pruned.
Our work focuses on modifying the Fisher-weighted merging method devised by Matena and Raffel (2022) with the insights gained from Kwon et al. (2022): that the Fisher information of masks can act as an importance measure for their affected parameter blocks. We insert masks and calculate an approximate of the Fisher information of each mask. By associating the mask Fisher information to relevant parameters in an empirical scheme, we merge models via weighted averaging similar to Matena and Raffel (2022). By only calculating the Fisher information of masks instead of the Fisher information of all parameters, our methods provide a massive reduction in computational cost while improving the performance benefits. As we restrict our context to performance on the respective tasks of fine-tuned model merges, our method does not require access to the validation set. The Fisher information matrix of each fine-tuned model on its respective task can provide a measure of how important each parameter is to that particular model, and we hypothesize that weighing by said Fisher information allows constructive merging, resulting in a single model performant in all given tasks.
We validate our hypothesis by demonstrating the effectiveness of our proposed methodology on various architectures of differing sizes, combining finetuned models on six tasks from the GLUE benchmark (Wang et al., 2018). Our method provides a significant improvement over simple averaging and Fisher-weighted merging in most models, using extremely small amounts of data and computational resources. We also note our method’s efficiency at scale, with an average speedup of 57.4x in comparison to Fisher-weighted merging on BERT (Large).
Our code is publicly released on Github. [^1] https://github.com/thennal10/fisher-nodes-merging
2.   Related Work
Much of the work in model merging has been done in addressing permutation symmetry when models are trained from different initializations (Ainsworth et al., 2023; Singh and Jaggi, 2023; Jin et al., 2023; Li et al., 2016; Tatro et al., 2020; Entezari et al., 2022). However, in the case of different fine-tuned models trained from the same parent pre-trained model, merging can often be done directly. As the optimization trajectory overlaps at the beginning, even simple averaging can provide reasonable results (Choshen et al., 2022; Wortsman et al., 2022; Ilharco et al., 2023; Jin et al., 2023). In this setting, model merging has been explored for a wide variety of contexts including improving single-task performance (Choshen et al., 2022; Wortsman et al., 2022), transfer learning (Matena and Raffel, 2022), federated learning (Li et al., 2020; McMahan et al., 2016), and improving generalization (Jin et al., 2023; Ilharco et al., 2023; Arpit et al., 2022; Cha et al., 2021).
Restricting ourselves to the specific challenge of merging fine-tuned task-specific models for multi-task learning, many methods improve over the baseline of simple parameter averaging (Jin et al., 2023; Yadav et al., 2023; Ilharco et al., 2023; Matena and Raffel, 2022). Fisher-weighted averaging, proposed by Matena and Raffel (2022), was originally evaluated in the context of transfer learning but it has been further employed for multi-task learning (Yadav et al., 2023).
The Fisher information matrix as a statistical measure has been used extensively in machine learning for a wide variety of applications (Kirkpatrick et al., 2017a; Kwon et al., 2022; Pascanu and Bengio, 2014; Soen and Sun, 2021; Hannun et al., 2021; Liu et al., 2021). Liu et al. (2021) prominently uses it for the task of pruning across multiple model architectures. Kwon et al. (2022) provides a fast Transformer-specific pruning scheme using the Fisher information matrix of mask nodes.
As the full Fisher information matrix is infeasible to calculate for most applications, only the diagonals are computed over a specified sample as an approximation (Matena and Raffel, 2022; Kirkpatrick et al., 2017a). However, even the diagonal approximation incurs the same computational cost as training on a specified number of examples. Our method borrows the mask architecture from Kwon et al. (2022), applying it to the field of model merging to reduce computational cost while keeping the performance benefits of Fisher-weighted averaging. We also note that our method does not require access to the validation set, in contrast to other methods (Matena and Raffel, 2022; Yadav et al., 2023; Ilharco et al., 2023).
3.   Methodology
Our approach relies on the hypothesis that the Fisher information of mask nodes can represent the importance of the parameters enclosed by the mask, as inspired by Kwon et al. (2022). Using the Fisher information of the mask nodes as a proxy for the Fisher information of the corresponding block of parameters, we merge parameters in a weighted-averaging merging scheme similar to Matena and Raffel (2022). We further consider only the training sets of the fine-tuned models for calculating the Fisher information in order to remove the dependency on the validation set and widen the applicability of our method.
In accordance with Kwon et al. (2022), our method focuses on the Transformer architecture, specifically the BERT architecture family (Devlin et al., 2019). BERT can be characterized as a stack of homogeneous Transformer encoder blocks, consisting of a multi-headed attention block followed by a feed-forward block. Masks are inserted on each head of the multi-headed attention block, and on each row in the intermediate linear layer of the feed-forward block, referred to as filters by Kwon et al. (2022).
Following the notation of the formal BERT algorithm given by Phuong and Hutter (2022), we represent the modified multi-headed attention algorithm in Equation 1 and the modified feed-forward layer in Equation 2:
For $\text{For }h\in\{1,...,H\}:$ (1)
$\quad Y_{h}=\texttt{Attention}(X;W^{h,l}_{qkv})$
$Y=\{Y_{1},Y_{2},...,Y_{H}\}$
$X\leftarrow X+W^{o}(Y\odot m^{l}_{mha})+b^{o}$
$\begin{split}X\leftarrow X+W^{l}_{mlp2}\texttt{GELU}(m^{l}_{mlp}\odot W^{l}_{mlp1}X+b^{l}_{mlp1})\\ +b^{l}_{mlp2}\end{split}$ (2)
We assume a given input $X$ and present only the modified section. We also omit the transposition of the bias vectors present in the original algorithm for conciseness.
In these equations, $m^{l}_{mlp}$ and $m^{l}_{mha}$ are mask vectors for the $l$-th layer of the feed-forward network and the multi-headed attention, respectively. $m^{l}_{mlp}=\mathbf{1}_{D\times 1}$ where $D$ represents the output size of the intermediate layer (the number of rows of $W^{l}_{mlp1}$, or the number of filters as per Kwon et al. (2022)), and $m^{l}_{mha}=\mathbf{1}_{H\times 1}$ where $H$ is the number of attention heads. The concatenated masks can thus be characterized as $m_{mha}=\mathbf{1}_{H\times L}$ and $m_{mlp}=\mathbf{1}_{R\times L}$, where $L$ is the total number of layers. As the masks are only used for their representational value, their values stay constant and equal to $1$ throughout the process.
The Fisher information matrix can be formulated in terms of the partial differentials of the loss function (Kwon et al., 2022):
$I:=\frac{1}{|D|}\sum_{(x,y)\in D}\left(\frac{\partial}{\partial m}L(x,y;1)\right)\left(\frac{\partial}{\partial m}L(x,y;1)\right)^{T}$ (3)
As the full Fisher information matrix takes $O(|\theta|^{2})$ to store, it becomes impractical for even a moderate number of masks. Thus, we use the diagonal approximation of the full matrix, as commonly practiced (Kirkpatrick et al., 2017b; Matena and Raffel, 2022; Kwon et al., 2022). The masks then have an associated Fisher information value equivalent to the mean squared sum of the gradients:
$I_{ii}:=\frac{1}{|D|}\sum_{(x,y)\in D}\left(\frac{\partial}{\partial m_{i}}L(x,y;1)\right)^{2}$ (4)
Matena and Raffel (2022) provide the following closed-form solution for model merging using the diagonal approximation of the Fisher information $F_{j}$ for parameters $\theta_{j}$ of the $j^{th}$ model, with $M$ as the total number of models to be merged:
$\theta^{*}=\frac{\sum_{j=1}^{M}\lambda_{j}F_{j}\theta_{j}}{\sum_{j=1}^{M}\lambda_{j}F_{j}}$ (5)
We keep $\lambda_{j}=1$ as we do not conduct hyperparameter tuning on the validation set. We approximate $F_{j}$ of each parameter with the Fisher information $I_{ii}$ of the mask node $m_{i}$ according to the following scheme:
  1. 1.
    If the parameter is in the query and key matrices of a particular attention head, i.e. $\theta_{j}$ is in $W_{qk}^{l,h}$ as per Equation 1, the Fisher information of the corresponding mask $m_{mha}^{l,h}$ is taken for $F_{j}$. Including the value parameters here was found to empirically lead to a lower performance.
  2. 2.
    If the parameter in a particular row $r$ of the intermediate layer of the feed-forward block, or $W^{l,r*}_{mlp1}$ as per Equation 2, the Fisher information of the corresponding mask $m^{l,r}_{mlp}$ is taken taken for $F_{j}$.
  3. 3.
    For all other parameters, $F_{j}=1$, resulting in merging equivalent to plain averaging.
We note that our methodology requires the calculation of gradients for only $(H+D)\times L$ parameters, in comparison to Fisher-weighted averaging requiring $|\theta|$, almost certain to be magnitudes higher.
4.   Evaluation
Model (normal-$\downarrow$) Averaging Fisher (2022) Ours
No. of Samples (normal-$\rightarrow$) 128 2,048 32,768 128 2,048 32,768
BERT (Base) 91.2 92.0 91.7 91.9 95.2 (+4.0) 92.8 (+1.6) 93.7 (+2.5)
BERT (Large) 85.8 87.0 86.1 89.2 90.1 (+4.3) 88.1 (+2.3) 87.9 (+2.1)
BERT (Tiny) 90.8 89.6 89.2 90.4 86.1 (-4.7) 88.4 (-2.4) 88.7 (-2.1)
RoBERTa 86.2 91.5 88.5 88.5 92.7 (+6.5) 89.4 (+3.2) 88.8 (+2.6)
Table 1: Normalized and aggregated metrics across tasks.
We evaluate our methodology on multiple pre-trained sizes of BERT, specifically tiny (Turc et al., 2019), base, and large (Devlin et al., 2019), along with the base model for the BERT derivative RoBERTa (Liu et al., 2019). We note however that with minimal modification it can be applied to all Transformer-based models.
[…]
For each checkpoint, we take 128, 2,048, and 32,768 samples from the training set of the corresponding dataset to calculate the Fisher information. We then merge the models, two at a time, and evaluate on the validation set of both the corresponding datasets. As the final classification head layer for each finetuned model is different and task-specific, we use the respective head for each task when evaluating. For comparison, the same procedure is conducted using the original Fisher-weighted merging methodology in (Matena and Raffel, 2022), along with simple averaging (Wortsman et al., 2022; Choshen et al., 2022).
5.   Results
Table 1 showcases the performance of our proposed method compared to simple averaging and the Fisher-weighted model merging of Matena and Raffel (2022). For ease of comparison, the difference between our method and the simple averaging is also indicated. Results are normalized based on the respective fine-tuned model’s performance on each task, and the median is taken to reduce the effects of outliers. The results demonstrate notable variations based on the model size and the number of samples used for calculating Fisher information.
For BERT (Base), our method demonstrates a consistent improvement over other methods across all sample sizes, with the most significant gain seen at 128 samples, providing a +4.0 increase in performance from simple averaging. BERT (Large) and RoBERTa follow a similar trend, with our method consistently outperforming the others. The greatest increase is again at 128 samples, with a performance boost of +4.0 and +6.5 respectively.
The results for BERT (Tiny) show an anomalous decrease in performance from simple averaging, in both Fisher-weighted merging and our method. However, we note that our method closes the gap as the number of samples increases, with a -2.1 decrease using 32,768 samples. We also stress the fact that BERT (Tiny) is an exceptionally small model, with only two transformer layers and a hidden embedding size of 128.
We note that our method also provides a significant efficiency advantage over Fisher-weighted averaging, with a 57.4x elapsed time speed-up for BERT (Large). To reduce the effect of overhead, the speedup is calculated using only the average runtime of the Fisher information approximation algorithm for each method, with an equal number of samples, running on a Nvidia T4 GPU.
Overall, these results confirm the efficacy of using the Fisher information of mask nodes for model merging, particularly in scenarios with limited data or computational resources. Further investigation is needed to fully understand the dynamics with scale, particularly for extremely small models like BERT (Tiny).
6.   Conclusion
In this paper, we introduced a novel model merging method for Transformers that builds upon the Fisher-weighted averaging methodology. By calculating the Fisher information of masks in attention heads and feed-forward layers, our approach significantly reduces the computational cost associated with full-scale Fisher-weighted merging while improving performance. Without a dependence on the validation set, our method is widely applicable for model merging in the context of multi-task learning.
We evaluated our method across several models and tasks, demonstrating its effectiveness and efficiency. The low computation cost enables practical model merging for real-world applications, particularly in resource-constrained environments. Future work could be done to evaluate our method’s capabilities for merging 3+ models at a time. Novel mask schemes may also be considered, along with an extension to different architectures.
7.   Bibliographical References
8.   Language Resource References
[…]
Appendix
1.   Checkpoints
For all architectures we mention, 6 models finetuned on the 6 respective tasks from the GLUE dataset were used. The checkpoints are given below.
1.1.   BERT Base
  • MNLI: JeremiahZ/bert-base-uncased-mnli
  • QQP: textattack/bert-base-uncased-QQP
  • QNLI: textattack/bert-base-uncased-QNLI
  • SST2: textattack/bert-base-uncased-SST-2
  • MRPC: textattack/bert-base-uncased-MRPC
  • RTE: textattack/bert-base-uncased-RTE
1.2.   BERT Large
  • MNLI: yoshitomo-matsubara/bert-large-uncased-mnli
  • QQP: yoshitomo-matsubara/bert-large-uncased-qqp
  • QNLI: yoshitomo-matsubara/bert-large-uncased-qnli
  • SST2: yoshitomo-matsubara/bert-large-uncased-sst2
  • MRPC: yoshitomo-matsubara/bert-large-uncased-mrpc
  • RTE: yoshitomo-matsubara/bert-large-uncased-rte
1.3.   BERT Tiny
  • MNLI: M-FAC/bert-tiny-finetuned-mnli
  • QQP: M-FAC/bert-tiny-finetuned-qqp
  • QNLI: M-FAC/bert-tiny-finetuned-qnli
  • SST2: M-FAC/bert-tiny-finetuned-sst2
  • MRPC: M-FAC/bert-tiny-finetuned-mrpc
  • RTE: muhtasham/bert-tiny-finetuned-glue-rte
1.4.   RoBERTa
  • MNLI: JeremiahZ/roberta-base-mnli
  • QQP: JeremiahZ/roberta-base-qqp
  • QNLI: JeremiahZ/roberta-base-qnli
  • SST2: JeremiahZ/roberta-base-sst2
  • MRPC: JeremiahZ/roberta-base-mrpc
  • RTE: JeremiahZ/roberta-base-rte